### v2

# EvoC2Rust: A Skeleton-guided Framework for Project-Level C-to-Rust Translation

Abstract.
Translating legacy C codebases to Rust is increasingly demanded for building safety-critical systems. While various approaches have emerged for this task, they face inherent trade-offs: rule-based methods often struggle to satisfy code safety and idiomaticity requirements, while LLM-based methods frequently fail to generate semantically equivalent Rust code, due to the heavy dependencies of modules across the entire codebase. Recent studies have revealed that both solutions are limited to small-scale programs. In this paper, we propose EvoC2Rust, an automated framework for converting complete C projects to equivalent Rust ones. EvoC2Rust employs a skeleton-guided translation strategy for project-level translation. The pipeline consists of three stages: 1) it first decomposes the C project into functional modules, employs a feature-mapping-enhanced LLM to transform definitions and macros, and generates type-checked function stubs, which form a compilable Rust skeleton; 2) it then incrementally translates functions, replacing the corresponding stub placeholders; 3) finally, it repairs compilation errors by integrating LLM and static analysis. Through evolutionary augmentation, EvoC2Rust combines the advantages of both rule-based and LLM-based solutions. Our evaluation on open-source benchmarks and six industrial projects demonstrates the superior performance of EvoC2Rust in project-level C-to-Rust translation. The results show that our approach outperforms the strongest LLM-based baseline by 17.24% in syntax accuracy and 14.32% in semantic accuracy, while also achieving a 43.59% higher code safety rate than the best rule-based tool.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the IEEE/ACM International Conference on Software Engineering: Software Engineering in Practice; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2025/02
1. Introduction
The translation of legacy C code to Rust is gaining significant attention in software engineering, spurred by critical memory safety concerns (Li et al., 2025; DARPA-SN-24-89, 2024). Such vulnerabilities constitute the most prevalent category of critical security defects in production C systems, accounting for approximately 70% of high-severity vulnerabilities in industry reports from Google and Microsoft (Lord, 2023). This has accelerated a paradigm shift toward compile-time safety guarantees through safe programming languages (Office of the National Cyber Director, 2024). In response, Rust has emerged as a leading alternative that offers memory safety by enforcing a strict ownership and borrowing model (Williams, 2024).
Figure 1. Overview of EvoC2Rust.
Recent work proposes rule-based tools (c2r, 2025; Zhang et al., 2023; Hong and Ryu, 2023), which convert C code to Rust using primarily syntax-based transformations on a terse intermediate representation of compilation. Such translation inherently generates non-idiomatic Rust code with low-level constructs and may also contain semantic inaccuracies. (Pan et al., 2024). Another line of work proposes LLM-based techniques (Eniser et al., 2024; Yang et al., 2024b; Nitin and Ray, 2024) to utilize the code comprehension capability of code LLMs. Due to the lack of parallel C-Rust training data and significant linguistic differences, these approaches often fail to produce semantically equivalent Rust code (Ou et al., 2024).
While automated code translation has been extensively investigated (Rozière et al., 2020; Zhu et al., 2022; Yang et al., 2024a), the translation of C-to-Rust projects poses special challenges:
Challenge-1: Substantial linguistic discrepancies for ensuring safety. C and Rust differ fundamentally in their safety models. C permits loose type checking, unrestricted pointer arithmetic, and manual memory management, whereas Rust enforces strict compile-time type safety, ownership-based memory guarantees, and disciplined reference usage through its borrowing rules. Existing rule-based translators often compromise safety by relying on raw pointers, unsafe blocks, or external C functions (Pan et al., 2024). While LLM-based approaches can generate code with improved safety, they often lack sufficient context understanding (e.g., variable scope or pointer lifetimes), potentially leading to logical errors or memory safety violations in the translated Rust code (Sim et al., 2025). As a result, neither rule-based nor LLM-based methods alone reliably meet Rust’s stringent safety requirements.
Challenge-2: Project-level code dependence. Translating complete C projects poses greater challenges than translating isolated functions because the translated Rust project must preserve cross-module dependencies and consistent project-level structure (Shiraishi and Shinagawa, 2024; Shetty et al., 2024). Code LLMs often struggle with such large-scale context, leading to broken references, inconsistent APIs, or incorrect module boundaries (Pan et al., 2024). Consequently, existing approaches tend to perform well only on small-scale programs (e.g., $\leq$100 LoC) (Emre et al., 2021; Eniser et al., 2024), and fall short in real-world project settings.
In this work, we propose EvoC2Rust, an automated framework for translating complete C projects into equivalent Rust code. To address the linguistic gap and ensure safety (Challenges 1), EvoC2Rust enhances LLMs with safety-preserving mappings across seven core linguistic features, including types, macros, functions, operators, syntax structures, globals, and variadic arguments. To enable project-level translation (Challenge 2), EvoC2Rust introduces a skeleton-guided translation strategy that proceeds in three stages. First, it analyzes the input C project to extract high-level metadata and decomposes the project into functional modules, generating a compilable Rust skeleton with placeholder functions. Next, it incrementally translates these stubs using LLMs augmented with the predefined feature mappings. Finally, it combines LLMs with static analysis to iteratively repair compilation errors and refine the output. This evolutionary process effectively blends the strengths of rule-based and learning-based approaches, striking a practical balance between correctness and safety in automated C-to-Rust migration.
We evaluate EvoC2Rust using an open-source benchmark Vivo-Bench (Vivo, 2025) and a self-collected dataset of six industrial projects. Experimental results demonstrate that EvoC2Rust significantly outperforms existing baselines in project-level C-to-Rust translation. In comparison with LLM-based approaches, EvoC2Rust improves the compilation pass rate by 17.24–50.84% and the line acceptance rate by 14.32–44.79%, while also achieving marginally better code safety. When compared to rule-based methods, EvoC2Rust attains a 43.59–96.79% higher code safety rate, while maintaining competitive performance in terms of compilation and line acceptance rates. Furthermore, at the module level, EvoC2Rust achieves a 92.25% compilation success rate and an 89.53% test pass rate on industrial projects.
The contributions of our work are summarized as follows:
  • We propose a novel skeleton-guided framework that overcomes the barrier of inter-module dependencies to enable the translation of complete C projects into Rust.
  • We define safety-preserving mappings between core linguistic features of C and Rust to enhance LLMs’ safety awareness.
  • We evaluate EvoC2Rust on both open-source and industrial scenarios. Experimental results show that EvoC2Rust outperforms existing baselines, effectively translating C projects to correct and safe Rust code.
Our source code and experimental data are publicly available at https://github.com/bbzswcf/EvoC2rust.
2. Approach
Figure 1 illustrates the architecture of EvoC2Rust. It contains three key steps: (1) Skeleton Construction decomposes the C project into modules and constructs a Rust project skeleton to assemble the Rust modules (Section 2.1); (2) Module Translation incrementally converts definitions, macros, and functions using safety-preserving linguistic mappings (Section 2.2); and (3) Post-Generation Repair combines LLM-based refinement with static analysis for code optimization (Section 2.3).
2.1. Project Skeleton Construction
A straightforward idea of project-level code translation is to translate each function individually and assemble them into a project. However, naively aggregating translated functions often leads to cascading interdependency errors. To address this, EvoC2Rust employs a skeleton-guided translation strategy, which first constructs a compilable Rust project skeleton based on project metadata and then incrementally populates it with translated function implementations.
Given a C project, EvoC2Rust first parses the source code using Tree-sitter (tre, 2025) to extract key syntactic elements and abstract them into a unified project metadata representation. This metadata captures structural information including filenames, include statements, macro definitions, type declarations, global identifiers, function signatures, and a mapping table between identifier declarations to source files. They are then systematically transformed into Rust equivalents via rule-based conversion. For instance, C include dependencies are transformed to Rust use imports, and identifier declarations are converted into Rust pub use statements (e.g., pub use {Rust_filename}::{identifier_name}) using the declaration-to-file table.
Using the converted metadata, EvoC2Rust constructs a Rust project skeleton—comprising Rust constructs automatically translated from C definitions, macros, and function signatures—that accurately preserves the structural organization of the original C project (see Figure 2). Each C function is assigned a type-checked dummy counterpart in Rust using its translated signature and the unimplemented!() macro. These placeholders serve as safe stand-ins, allowing the project to compile successfully even before the actual function bodies are translated.
The skeleton provides a stable framework that allows function bodies to be translated and integrated incrementally. This ensures each addition is contextually correct and free from dependency errors, enabling a seamless and scalable translation process even for large-scale projects with complex interactions.
Figure 2. An illustration of translating a simple C project to Rust.
Table 1. Type Mappings between C and Rust
Type C Type C Example Rust Type Rust Example
Array int[N] int a[3] = {1, 2, 3} Array<i32, N> let mut a: Array<i32,3> = arr![1,2,3]
Pointer int * int *a = &b Ptr<i32> let mut a: Ptr<i32> = c_ref!(b)
String char * char *c = "Hello World!" Ptr<u8> let mut c=cstr!("Hello World!")
Function Pointer (*func) typedef int (*MyFunc) (const void *, const void *) FuncPtr pub type MyFunc=FuncPtr<fn(Ptr<Void>, Ptr<Void>) -> i32>
File FILE * FILE *f = fopen(name, "rb") FilePtr let mut f=c_fopen!(name, cstr("rb"))
Variadic Argument va_list va_start(arg_ptr, prev) VaList (…, prev: …, arg_ptr: VaList)
2.2. Module Translation with Safety-preserving Mapping
EvoC2Rust employs an LLM to automatically translate C modules into Rust constructs, converting definitions, macros, and function signatures to assemble the skeleton, and then transforming function bodies to populate the placeholder dummy functions within the skeleton. To preserve safety and semantic alignment during translation, EvoC2Rust augments the LLM with a comprehensive set of linguistic mappings across seven key categories:
1) Type Mapping. Each C type is associated with a semantically consistent Rust counterpart, as summarized in Table 1. For example, to represent C-style strings, we introduce a cstr!() macro in Rust that guarantees null termination, thereby avoiding inconsistencies in downstream operations such as encryption and compression. Beyond simple type correspondences, we also capture the operational semantics of C types. Since C pointers permit arithmetic (e.g., addition, subtraction), the address-of operator, and array-style subscripting, we implement equivalent traits in Rust to mirror these behaviors. Likewise, to model the implicit decay of C arrays into pointers, we provide an explicit cast() method. Together, these mappings reduce the complexity of translating raw C pointers by supplying direct Rust equivalents for a wide range of pointer classes and operations.
2) Type Conversion. We introduce a custom CastIntoTyped trait to handle type conversions in Rust, supporting both integer-to-integer and pointer-to-pointer conversions. For safety, pointer-to-integer casts are allowed, but integer-to-pointer casts are strictly disallowed. For instance, the C expression ‘a = (uint16_t) b’ is translated as ‘a = b.cast::<u16>()’, and ‘pi = (int*)pu’ becomes ‘pi = pu.cast::<Ptr<i32>>()’. To simulate C’s implicit conversion semantics in Rust’s type-inferred context, we also provide a typeless cast() method. This proves particularly effective during the repair phase, where it substantially increases the success rate of LLM-driven optimization by efficiently resolving type errors stemming from implicit conversions in the original C code.
3) Macro/Function Mapping. We provide Rust counterparts for native C macros such as NULL and LINE, as well as for common C library functions including malloc, free, and memcpy. These implementations prioritize safe abstractions over direct invocations of unsafe primitives. For example, rather than relying on the unsafe std::ptr::copy to implement memmove, we employ a safer alternative. By explicitly incorporating these mappings into the prompt, we guide the LLM to apply them consistently, enabling systematic translations such as converting the C function name() into the Rust macro c_name!().
4) Syntax Structure Mapping. Certain C control flow constructs lack direct semantic equivalents in Rust, posing a risk of logical errors if translated literally. Specifically, Rust’s for loop operates differently from C’s, C’s do while has no native counterpart, and switch statements are not fully aligned with Rust’s match. To ensure semantic fidelity, we implement dedicated macros—c_for, c_do, and c_switch—that replicate the exact behavior of their C counterparts. These are introduced to the LLM via targeted prompts, ensuring the original syntax is replaced with the appropriate macro calls.
5) Operator Mapping. We provide Rust equivalents for C-specific operators such as ++, --, &, and sizeof. The sizeof operator is mapped to two distinct macros: c_sizeof!() and c_sizeofval!(). The increment and decrement operators (++, --) are implemented as four generic functions to handle both prefix and postfix forms for all integer and pointer types. Notably, the address-of operator & is not translated to Rust’s reference operators (& or &mut), but to a dedicated c_ref!() macro that correctly obtains a pointer to the value.
6) Global Variable Mapping. Because mutable global variables (static mut) are inherently unsafe in Rust, we introduce a Global<T> wrapper type to manage them safely. This wrapper employs an internal Mutex to ensure thread safety, thereby avoiding the need for unsafe code blocks. To further simplify usage, we provide a global!() macro that supports lazy initialization of global variables.
7) Variadic Argument Mapping. Since variadic arguments in C are most commonly used for logging and string formatting, we model the VaList type in Rust as a Slice of references to values implementing the Display trait. To support formatting operations, we provide a va_format!() macro, which in turn enables Rust implementations of variadic C functions such as snprintf.
Table 2. An Example of Transformation Patterns
Pattern Name: Variadic Argument Mapping   Motivation: Rust does not support C-style variadic functions (e.g., va_list).   Consequence: Current LLMs cannot directly translate C’s variadic features.   Solution: Define a VaList type and va_format! macro: pub type VaList<‘a> = &‘a [&‘a dyn Display]; macro_rules! va_format { … }   Example: C Code: int snprintf(char *str, size_t size, const char *format, …) { … } Rust Code: pub fn snprintf(mut buf: Ptr<u8>, size: usize, format: Ptr<u8>, va: VaList) -> i32 { let mut fmt = format.to_string(); … return length; }
These feature mappings are formulated as transformation patterns, each specifying the motivation, consequence, solution, and corresponding code examples, as illustrated in Table 2. By distilling expertise from senior practitioners, these patterns encode tacit professional knowledge into explicitly reusable C-to-Rust migration guidelines.
To steer the LLM toward accurate and idiomatic translations, we retrieve the most relevant patterns based on the input C code. Specifically, we encode both the code snippet and pattern examples into dense vectors, compute cosine similarity, and select the top-K matching patterns. These retrieved patterns are then injected into the LLM prompt using the following structured template.
Prompt Template for Translating C Modules to Rust Translate the C [macro/definition/function] to Rust.
Patterns: {retrieved transformation patterns}
Demonstrations: {paired C-to-Rust examples}
C Source Code: {input source code}
2.3. Post-Generation Repair
Finally, EvoC2Rust refines the translated Rust code through a compilation-driven, cascading repair process that combines syntactic rules with LLM-based correction. For each problematic snippet, it generates multiple repair candidates based on compiler feedback, retaining only those that reduce error counts until no further improvements are possible.
While rule-based techniques efficiently resolve trivial and well-defined errors, LLMs excel at addressing complex semantic issues. A hybrid strategy—applying rules first, followed by LLM refinement—yields the best results. We observe, however, that persistent low-level syntax errors such as bracket mismatches significantly impede both repair methods, particularly rule-based correction. To mitigate this, our approach prioritizes bracket repair before other fixes, ensuring that basic syntax errors are resolved prior to semantic analysis. This allows the Rust compiler to provide clearer diagnostic information for subsequent repair stages.
Specifically, EvoC2Rust implements a three-step repair chain:
1) Bracket Repair: An LLM analyzes compiler outputs and error messages to correct mismatched brackets and similar syntax issues. The prompt includes instructions, exemplar fault-correction pairs, the erroneous Rust snippet, and corresponding compilation errors.
Prompt Template for Bracket Repair Fix the compilation bugs in the following Rust code with the provided compilation error messages, possibly because of mismatched parentheses. Only correct lines that have unmatched parentheses bugs; do not modify any other code.
Demonstrations: {example pairs of incorrect/correct Rust code}
Rust Source Code: {input source code}
Compilation Error Messages: {input error messages}
2) Rule-Based Repair: We define a set of syntax repair rules as regex-based transformation patterns to address common syntax-level issues. These include adjusting misplaced derive macros, removing redundant cast() operations, and resolving problematic array access patterns such as s[s.i], which can cause simultaneous mutable and immutable borrow conflicts.
3) LLM Refinement: In the final step, the LLM addresses residual semantic discrepancies and intricate compilation errors, such as type inconsistencies, unidiomatic usage, or structural mismatches. The prompt incorporates guidelines, fault-correction examples, the erroneous Rust translation, and relevant compilation feedback. These exemplar pairs, which are manually crafted by Rust experts and refined through iterative experimentation, cover common recurring errors to enable the LLM to generalize effective correction strategies (see our repository for detailed examples).
Prompt Template for LLM Refinement Fix the compilation bugs in the following Rust code according to the compilation information.
Instruction: {refinement guidelines}
Demonstrations: {example pairs of incorrect/correct Rust code}
Rust Source Code: {input source code}
Compilation Information: {input compilation feedback}
3. Experimental Setup
We conduct experiments to evaluate the effectiveness of EvoC2Rust, aiming to answer the following research questions.
  • RQ1: How effectively does our method translate complete C projects to equivalent safe Rust code?
  • RQ2: What is the module conversion accuracy achieved by our method?
  • RQ3: To what extent do key components contribute to the overall performance of our method?
  • RQ4: How does our method scale to large projects in terms of accuracy, safety, and efficiency?
3.1. Comparison Methods
We compare EvoC2Rust against three categories of C-to-Rust translation methods: rule-based (C2Rust), LLM-based (Self-Repair and Tymcrat), and hybrid techniques (C2SaferRust). We also incorporate a direct LLM prompting baseline that translates code without repository context.
Specifically, the baseline methods evaluated in our study are as follows:
  • C2Rust (c2r, 2025): a rule-based C-to-Rust translator that converts C code to Rust via AST analysis and manually defined transformation rules.
  • C2SaferRust (Nitin et al., 2025): a hybrid technique built on C2Rust, augmented with LLM-based post-processing and test-driven optimization to enhance the safety and idiomaticity of the translated Rust code. We follow the original configuration and use 5 optimization iterations.
  • Self-Repair (Khatry et al., 2025): a multi-agent self-repair framework for project-level C-to-Rust translation. It incrementally performs file translation and uses iterative compilation and testing feedback to correct errors automatically. We adopt the default setting of 3 self-repair rounds per task.
  • Tymcrat (Hong and Ryu, 2025b): a project-level translation method that improves translation quality through type migration. It generates multiple candidate Rust signatures and refines them iteratively using compiler feedback. We use the default configuration, which generates 4 candidate signatures per function.
  • Direct Prompting: a baseline that applies zero-shot LLM prompting for direct C-to-Rust translation. To address its limitations in preserving project-wide consistency and resolving dependencies, we integrate it within our project skeleton framework.
We implement all baseline methods using their officially released code, while substituting their underlying models with those used in our study to ensure a fair comparison.
Table 3. Statistics of the Datasets
Dataset Projects Files LoC Functions Macros (Functions) Definitions Test Cases
[…]
C2R-Bench avl 9 836 29 29 9 121
bzp 18 2258 92 82 14 17
md5 2 324 8 20 1 37
sha256 2 280 9 11 2 12
rapidlz 7 748 28 38 10 12
cmptlz 25 3724 122 166 27 23
Total 63 8170 288 346 63 222
3.2. Datasets
We evaluate EvoC2Rust using two benchmarks: the open-source Vivo-Bench (Vivo, 2025), and our newly introduced C2R-Bench, which comprises six industrial projects.
Vivo-Bench is derived from the 2025 Vivo C-to-Rust Innovation Competition and consists of 19 algorithmic C projects. These projects contain 200 functions, 95 definitions, and 29 non-function macros, with each project spanning 1–3 files (33–630 LoC per file). Two senior engineers produced the reference Rust translations with LLM assistance. Subsequently, they validated the translations by adding tests to achieve complete function coverage, resulting in a total of 113 top-level test cases.
To assess translation performance in industrial settings, we constructed C2R-Bench using six production projects from Huawei’s software ecosystem. These are single-threaded user-space applications using only standard C libraries, characterized by complex multi-file architectures (280–3,724 LoC per file) and extensive cross-file dependencies. Three senior engineers produced the reference Rust implementations via LLM-assisted translation, which were validated using 222 top-level test cases.
Each project in both Vivo-Bench (Vivo, 2025) and C2R-Bench provides four artifacts: the original C source code, C test cases, the reference Rust translation, and corresponding Rust test cases. Table 3 summarizes the dataset statistics.
3.3. Evaluation Strategy and Metrics
Our evaluation assesses C-to-Rust translation performance at two granularities: project-level and module-level, each employing a tailored strategy and specific metrics.
Project-level Evaluation. We benchmark the translation of complete software projects under realistic conditions, where reference implementations are unavailable. This strategy employs incremental compilation: we construct a skeleton with placeholder modules and iteratively replace them with translated code, reverting any modules that cause compilation failures. We employ three metrics to quantify project-level performance:
  • Incremental Compilation Pass Rate (ICompRate): The proportion of translated modules that compile successfully when integrated step-wise into the project skeleton, measuring syntactic correctness  (Shiraishi and Shinagawa, 2024).
  • Line Acceptance Rate (AccRate) (Ziegler et al., 2022): This metric evaluates the fidelity of the initial translation by comparing it to the manually corrected version. Precision captures the percentage of correct lines in the initial output, while Recall measures their preservation in the final code.
  • Code Safety Rate (SafeRate) (Ling et al., 2022): The proportion of generated Rust code that is memory-safe, i.e., free of unsafe functions or blocks.
Module-level Evaluation. For a finer-grained analysis, we assess translation accuracy at the module level using a fill-in-the-blank strategy. Given a C module and its reference Rust implementation, we remove the latter, translate the C module, and insert the result into the project to check compilation and test outcomes. Two metrics are used for this assessment:
  • Fill-in Compilation Pass Rate (FCompRate): The percentage of generated modules that compile correctly upon integration.
  • Test Pass Rate (TestRate) (Khatry et al., 2025): The percentage of generated modules that pass all associated unit tests.
3.4. Implementation Details
Table 4. Performance of Various Methods in Translating Complete C Projects to Rust
Dataset Method ICompRate AccRate SafeRate
Precision Recall
[…]
C2Rust (c2r, 2025) 100 100 100 0
C2SaferRust (Nitin et al., 2025) 100 100 100 60.00
[…]
Self-Repair (Khatry et al., 2025) 84.21 92.08 89.24 87.60
Tymcrat (Hong and Ryu, 2025b) 87.35 92.11 92.44 97.88
Direct Prompting 35.49 55.15 51.95 77.09
EvoC2Rust (Ours) 100 99.83 99.86 98.00
[…]
Self-Repair (Khatry et al., 2025) 52.63 33.12 38.36 77.57
Tymcrat (Hong and Ryu, 2025b) 81.76 76.82 74.47 87.11
Direct Prompting 26.23 21.72 27.51 89.81
EvoC2Rust (Ours) 87.65 87.92 83.45 98.22
[…]
C2Rust (c2r, 2025) 99.28 98.99 98.98 1.83
C2SaferRust (Nitin et al., 2025) 99.28 97.47 97.68 48.24
[…]
Self-Repair (Khatry et al., 2025) 49.21 72.35 13.79 83.92
Tymcrat (Hong and Ryu, 2025b) 72.02 77.82 74.94 95.92
Direct Prompting 56.67 55.65 52.70 82.43
EvoC2Rust (Ours) 93.84 97.56 97.34 97.41
[…]
Self-Repair (Khatry et al., 2025) 39.40 45.38 10.65 87.15
Tymcrat (Hong and Ryu, 2025b) 70.56 66.72 65.91 82.75
Direct Prompting 35.29 24.01 30.41 93.63
EvoC2Rust (Ours) 82.35 85.75 82.59 97.40
  • * Bold and underline denote the best and second-best results, respectively.
Table 5. Effectiveness of EvoC2Rust in Module-Level C-to-Rust Translation
Dataset Project FCompRate TestRate
[…]
Vivo-Bench 19 projects 99.07 98.50
C2R-Bench avl 100.00 92.53
bzp 95.21 92.55
md5 96.55 86.21
sha256 100.00 100.00
rapidlz 96.05 92.11
cmptlz 86.98 86.35
Subtotal 92.25 89.53
[…]
Vivo-Bench 19 projects 87.65 84.57
C2R-Bench avl 97.01 82.09
bzp 89.36 88.83
md5 89.66 89.21
sha256 90.91 90.91
rapidlz 78.95 77.63
cmptlz 70.79 68.89
Subtotal 80.63 77.91
We implement EvoC2Rust using DeepSeek-V3 (DeepSeek-AI, 2025) and Qwen3-32B (Alibaba, 2025) as the foundation models. Translation tasks use greedy decoding with max_tokens set to 4096. We leverage Tree-sitter v0.22.3 (tre, 2025) for C code parsing, the standard rustc (rus, 2025) for compilation, and the BGE-M3 model (of Artificial Intelligence, 2025) to generate embeddings and retrieve the top-10 most relevant patterns based on cosine similarity. Iteration limits are enforced to ensure practicality: 5 rounds for bracket repair and 3 rounds for LLM refinement.
Verified ground-truth implementations were established for the line acceptance rate metric as follows. The translation for each project was first executed by three senior engineers using the target methods. The generated Rust artifacts were then iteratively compiled, tested, and manually repaired with Claude Sonnet 4 (Anthropic, 2024) until all tests passed. A final code review by three authors confirmed the semantic equivalence of these implementations.
All experiments are conducted on a Linux server running Ubuntu 23.10, equipped with two NVIDIA 4090 GPUs and CUDA 12.4.
4. Results and Analysis
4.1. Main Results (RQ1)
EvoC2Rust demonstrates superior performance in project-level C-to-Rust translation, as summarized in Table 4. It consistently outperforms all baseline methods across both datasets, achieving the highest overall performance in terms of compilation success, line acceptance, and code safety.
EvoC2Rust demonstrates a significant advantage in generating memory-safe Rust code compared to rule-based and hybrid baselines. Although C2Rust and C2SaferRust achieve near-perfect compilation pass rates (e.g., 100% on Vivo-Bench), their safety performance is considerably lower. C2Rust exhibits code safety rates of only 0% and 1.83% on the two datasets, while C2SaferRust, despite LLM-enhanced optimization, reaches merely 60% and 48.24%. This confirms that syntactic translation, even with LLM post-processing, fails to reliably meet Rust’s safety requirements. In contrast, our method proactively prevents unsafe code by leveraging safety-guaranteed feature mappings.
Among LLM-based approaches, EvoC2Rust demonstrates consistent superiority. On Vivo-Bench, it surpasses all baselines by 12.65% to 64.51% in compilation success and 7.57% to 46.30% in line acceptance using DeepSeek-V3, while maintaining a 98% code safety rate. The performance gap widens on the complex C2R-Bench, where it achieves 93.84% compilation success (21.82%-44.63% higher than baselines) and over 97% line acceptance. This trend is also observed with the Qwen3-32B model, highlighting the effectiveness of our skeleton-guided, feature mapping-augmented translation paradigm.
Specifically, EvoC2Rust improves compilation success by 44.63% and line acceptance by 54.38% over Self-Repair on C2R-Bench with DeepSeek-V3 (42.95%/55.66% with Qwen3-32B). Self-Repair’s file-level processing often exceeds LLM context limits, leading to code truncation. Compared to Tymcrat, EvoC2Rust achieves a 21.82% higher compilation rate on C2R-Bench with DeepSeek-V3 (11.79% with Qwen3-32B). While Tymcrat employs diverse signature generation and context enrichment to reduce initial errors, its lack of expert-derived structured translation guidance ultimately constrains its effectiveness in complex scenarios. Direct Prompting performs poorest, with a 38.42% average compilation success and low line acceptance (39.13% precision, 40.64% recall). This is primarily due to the inherent linguistic gaps between C and Rust. The use of our skeleton framework, which helps manage project-wide context, prevents these metrics from being even lower.
Answer to RQ1. EvoC2Rust surpasses all baseline methods in project-level C-to-Rust translation, demonstrating the highest overall performance across compilation success, line acceptance, and code safety on both open-source and industrial benchmarks.
Table 6. Ablation Results on Key Components of EvoC2Rust
Variant Project Translation Module Translation
ICompRate SafeRate FCompRate TestRate
EvoC2Rust (Ours) 93.84 97.41 92.25 89.53
-w/o Repair#3 89.57 ( -4.27) 97.40 ( -0.01) 87.23 ( -5.02) 85.37 ( -4.16)
-w/o Repair#2-3 75.24 (-18.60) 97.46 ( +0.05) 82.21 (-10.04) 80.92 ( -8.61)
-w/o Repair#1-3 74.29 (-19.55) 97.21 ( -0.20) 81.06 (-11.19) 79.91 ( -9.62)
-w/o Feature Mapping and Repair#1-3 56.67 (-37.17) 82.43 (-14.98) 33.00 (-59.25) 30.27 (-59.26)
  • * Repair#1: bracket repair; Repair#2: rule-based repair; Repair#3: LLM refinement.
4.2. Module Translation (RQ2)
We further evaluate the module translation accuracy of EvoC2Rust, which serves as its core technical component. As shown in Table 3.4, EvoC2Rust exhibits strong performance at the module level. Using DeepSeek-V3, it achieves 100% compilation and 99.07% test pass rates on Vivo-Bench, while maintaining robust results on C2R-Bench (92.25% compilation, 89.53% test pass). Even with the smaller Qwen3-32B model which has 20× fewer parameters, EvoC2Rust still delivers competitive performance (80.63% compilation, 77.91% test pass), demonstrating its model-agnostic applicability.
A notable observation is the strong correlation between test pass rates and compilation success: modules that compile correctly generally also pass functional tests. This consistency results from our method’s dual mechanisms. Feature mapping prevents semantic errors, and compiler-guided repair maintains logically sound error correction.
Our analysis confirms that project complexity significantly impacts translation quality. Complex projects with intricate control flows and data dependencies consistently challenge LLM comprehension capabilities. For instance, the cmptlz project, featuring complex macros and dependencies, achieves test pass rates of 86.35% and 70.79% across models. In contrast, the simpler sha256 project attains optimal 100% pass rates.
Answer to RQ2. Our method demonstrates robust performance in module-level translation, achieving 92.25%–100% compilation success and 77.91%–99.07% test pass rates across model scales, confirming its effectiveness in handling projects of varying complexity.
4.3. Ablation Study (RQ3)
To assess the contribution of each component in EvoC2Rust and validate our design choices, we perform an ablation study by incrementally removing its key components: feature mapping, bracket repair, rule-based repair, and LLM refinement. Each variant is evaluated on C2R-Bench benchmark using the DeepSeek-V3 model. In our cascaded repair chain, later repairs depend on the output of earlier steps. Removing an early component while keeping later ones would thus propagate unresolved errors, preventing compilation or testing and invalidating the assessment of downstream components. To avoid this issue and ensure a meaningful evaluation, we remove components in reverse order, beginning with the final step (LLM refinement).
As illustrated in Table 6, every component in EvoC2Rust contributes critically to its overall effectiveness. Among them, the feature mapping mechanism is the most impactful component. By ablating this mechanism, both ICompRate and FCompRate drop sharply from 74.29% to 56.67% and from 81.06% to 33%, respectively. The impact on semantic correctness is also substantial, with TestRate decreasing from 79.91% to 30.27%. This result highlights the mechanism’s essential role in bridging C and Rust constructs through predefined transformation patterns that map C idioms to safe Rust equivalents where direct counterparts are absent.
The cascading repair chain also facilitates C-to-Rust translation. Removing all three repairs (-w/o repair#1-3) results in considerable performance degradation, with ICompRate dropping by 19.55%, FCompRate by 11.19%, and TestRate by 9.62%. Within this chain, rule-based repair (repair#2) contributes most significantly, raising the compilation pass rate by 9.68% and the test pass rate by 4.45% on average. LLM refinement (repair#3) further enhances syntactic correctness by 4.27% and semantic correctness by 8.98%, without compromising code safety.
Code safety remains consistently high (around 97%) across nearly all ablation settings. This stability is attributable to the feature mapping mechanism’s encapsulation of low-level operations within verified safe constructs. Only when both feature mapping and repairs are disabled does safety drop sharply to 82.43%, demonstrating the necessity of our integrated design.
Answer to RQ3. Each component in EvoC2Rust plays a crucial role in its overall performance. The feature mapping mechanism provides a safe and reliable basis by bridging C-to-Rust semantic gaps, while the cascading repair chain fixes remaining errors to maximize syntactic and semantic correctness.
4.4. Scalability on Larger Projects (RQ4)
To evaluate the scalability and performance of our method on large-scale codebases, we conduct experiments on 10 larger projects (9,973 - 91,588 LoC) from RepoTransBench (Wang et al., 2024), a benchmark derived from highly-starred GitHub repositories. During module translation, we exclude functions that depend on third-party libraries, as our current implementation primarily supports dependencies on the ISO C standard library. Using DeepSeek-V3 as the backbone model, EvoC2Rust achieves acceptable performance, with an ICompRate of 75.61% and a SafeRate of 96.20%. We do not report line acceptance rates due to the prohibitive cost of manually creating verified Rust references for projects at this scale.
By integrating results from all evaluated datasets (i.e., Vivo-Bench, C2R-Bench, and RepoTransBench), we further analyze translation performance across different project scales.
Translation Accuracy vs. Project Scale. As shown in Figure 3, compilation success (ICompRate) exhibits a clear negative correlation with project size. While the rate declines gradually from 100% for small projects ($\leq$500 LoC) to approximately 69% for large projects (¿10,000 LoC), it does not collapse. This robustness to scaling is due to our skeleton-guided approach, which effectively manages complex inter-module dependencies.
Translation Safety vs. Project Scale. The SafeRate remains consistently high (around 97%) across all project sizes, and even exhibits a slight improvement in larger projects. We attribute this stability to our safety-guaranteed feature mappings, which encapsulate low-level operations into a relatively fixed set of unsafe code blocks. As project size increases, the proportion of such unsafe code naturally decreases, resulting in the observed improvement in SafeRate.
Translation Efficiency vs. Project Scale. Figure 4 presents the translation efficiency, quantified as seconds per line of code (Sec/LoC), across projects of varying scales. The per-line translation time ranges from approximately 0.35 to 0.85 seconds, averaging 0.63 seconds. Notably, for projects exceeding 5,000 LoC, efficiency stabilizes at approximately 0.55 Sec/LoC, demonstrating strong scalability of our method. These performance fluctuations show no correlation with code length and are instead determined by project-specific characteristics. The primary factor influencing this variation is the number of LLM refinement iterations required, which depends directly on the inherent complexity of the code in each project.
Answer to RQ4. Our method exhibits strong scalability on large-scale projects, maintaining resilient compilation rates, high safety guarantees, and stable efficiency without performance collapse.
Figure 3. Translation performance at different project scales.
Figure 4. Translation time at different project scales.
4.5. Case Study
Figure 5. A case study on translating the rb_tree_rotate function.
We present a case study on the rb_tree_rotate function, a core operation in a red-black tree implementation, to illustrate the effectiveness of EvoC2Rust (Figure 5). This function involves intricate pointer manipulation, characteristic of low-level C code that is difficult to translate into safe, idiomatic Rust. Its complexity exposes common limitations of existing translation approaches.
As shown in Figure 5, baseline methods exhibit critical failures. Direct Prompting produces a project-level interface mismatch by passing a raw pointer (*mut) where a mutable reference (&mut) is required, revealing a lack of project-wide context awareness. Moreover, its heavy use of raw pointers necessitates unsafe blocks, contradicting Rust’s memory safety goals. Self-Repair attempts a more idiomatic translation using Box<T>, but introduces a function-level ownership error (use of moved value), demonstrating the difficulty of refactoring C-style memory management into safe Rust idioms without safety-guaranteed feature mappings.
In contrast, EvoC2Rust generates a correct and safe translation. This is achieved through our feature mapping mechanism, which employs a unified Ptr<T> smart pointer to preserve C-like pointer semantics while ensuring project-wide type consistency. Combined with a flexible .cast() method for type conversions, this approach resolves both the interface mismatch and ownership issues that cause other methods to fail. By abstracting low-level C operations into a safe Rust layer, EvoC2Rust succeeds where other techniques fall short.
5. Limitations and Threats to Validity
We have identified the following limitations and potential threats to the validity of our method:
Internal Validity. The main internal threat comes from correctness verification, which currently relies on predefined test cases. Adopting more robust validation techniques like fuzzing (Kadron et al., 2024) and self-debugging (Kang et al., 2025; Shi et al., 2024) could enhance test coverage and uncover subtle semantic errors. For line acceptance rate measurement, we use Claude-generated outputs verified by human experts as ground truth. While this manual validation ensures correctness beyond automated testing, it may not scale to larger datasets. Future work should explore semi-automated verification to balance accuracy and efficiency.
External Validity. Three primary threats affect external validity: 1) Generalizability: While our translation pipeline is designed to support multiple language pairs, the current implementation specifically targets C-to-Rust translation. Moreover, due to resource constraints, our evaluation is limited to DeepSeek-V3 and Qwen3-32B. These choices do not, however, affect the validity of our framework evaluation, as the primary contribution lies in the architecture itself rather than in model comparison. Extending the evaluation to include additional languages and models remains an important direction for future work. 2) Data leakage: Our method uses LLMs to translate C code to Rust. Since these models may have been pre-trained on public benchmarks, we mitigate this risk by including six industrial projects alongside open-source datasets in our evaluation. 3) Dataset characteristics: The current benchmarks are limited to single-threaded, user-level C projects with ISO C standard library dependencies. Future work should consider more complex settings such as multithreading, third-party libraries, and kernel-level code.
6. Related Work
6.1. C-to-Rust Translation
Existing work for automatic C-to-Rust translation falls into three broad categories: rule-based, LLM-based, and hybrid techniques.
Rule-based Methods rely on manually designed rules to transform C code into Rust. The C2Rust transpiler (c2r, 2025) exemplifies this approach by using Clang’s AST to generate Rust code, though the output relies heavily on unsafe blocks. Subsequent work has aimed to improve safety and idiomaticity: Emre et al. (Emre et al., 2021, 2023) used compiler feedback for pointer safety; Zhang et al. (Zhang et al., 2023) applied static ownership analysis; Ling et al. (Ling et al., 2022) and Hong et al. (Hong and Ryu, 2023, 2024a, 2024b, 2025a) focused on API safety and specific constructs; Larson et al. (Larsen, 2024) proposed pointer derivation graphs to translate unsafe Rust to safer Rust; Wu et al. (Wu and Demsky, 2025) retyped polymorphic C void pointers into uses of Rust generics. While effective, these methods require significant manual effort and often produce Rust code that is both unsafe and non-idiomatic.
LLM-based Methods leverage LLMs to generate idiomatic Rust code without relying on manually defined rules. Eniser et al. (Eniser et al., 2024) introduced FLOURINE, which uses fuzz testing to validate semantic equivalence between C and Rust code. Yang et al. (Yang et al., 2024b) adapted MSWasm-based testing for multi-language translation to Rust. Luo et al. (Luo et al., 2025) proposed IRENE, an LLM-based framework that integrates rule-augmented retrieval and structured summarization for function-level translation. Nitin et al. (Nitin and Ray, 2024) enhanced LLM translation by integrating static specifications and I/O tests. Xu et al. (Xu and Huang, 2025) optimized type migration by leveraging data flow graphs. Farrukh et al. (Farrukh et al., 2025) developed a multi-agent framework featuring iterative error correction. However, LLM-based translation suffers from lower accuracy due to the lack of parallel C-Rust training data (Ou et al., 2024).
Hybrid Methods seek to combine the strengths of both paradigms. For instance, C2SaferRust (Nitin et al., 2025) and PR2 (Gao et al., 2025) employ LLMs to refine rule-based translations, improving safety and idiomaticity while preserving functional equivalence through testing. SACTOR (Zhou et al., 2025) delegates different tasks to each paradigm, applying C2Rust for data types and an LLM for global variables and functions. Our method extends this direction with a two-stage framework: it first uses linguistic feature mappings (i.e., transformation rules) to guide LLM translation, then applies a compilation-driven repair mechanism combining LLM refinement with static analysis. It achieves an effective balance between translation accuracy and safety in automated C-to-Rust migration.
6.2. Project-level Code Translation with LLMs
While most C-to-Rust translation methods target small code units, recent work explores LLM-based project-level translation. These approaches typically decompose a C project into dependency-ordered units, translate them sequentially, and reassemble the results.
Shiraishi et al. (Shiraishi and Shinagawa, 2024) pioneered this direction by using project metadata to maintain consistency, though their method prioritizes compilation success over functional correctness. Syzygy (Shetty et al., 2024) and RustMap (Cai et al., 2025) employ a strategy of translating both functions and test cases while utilizing feedback mechanisms for repair. Nevertheless, these approaches necessitate complete dependency contexts (occasionally through manual intervention) and exhibit limited scalability when applied to complex industrial codebases. Khatry et al. (Khatry et al., 2025) and Ou et al. (Ou et al., 2024) introduced repository-scale benchmarks and self-repair techniques, yet rely on annotated Rust interfaces rarely available in practice. Hong et al. (Hong and Ryu, 2025b) addressed this with Tymcrat, a type inference-based method for automating signature translation.
Differing from existing approaches that translate modules sequentially following dependency order and require full project contexts, EvoC2Rust introduces a skeleton-guided strategy, which first constructs a compilable Rust skeleton and then executes incremental function translation. Beyond enabling concurrent processing, this design also enhances correctness by isolating the LLM from cross-module dependencies.
7. Conclusion
We present EvoC2Rust, a novel LLM-powered framework for automated translation of complete C projects to Rust. By leveraging feature mapping-enhanced LLMs, EvoC2Rust performs skeleton-guided code translation, augmented with hybrid error correction that combines LLM capabilities with static analysis. Extensive evaluation on both open-source and industrial benchmarks demonstrates that EvoC2Rust consistently achieves superior overall performance in syntax accuracy, semantic equivalence, and memory safety. Future work will extend the approach to more complex settings, including multithreading, third-party libraries, and kernel-level code across diverse application domains.